[Brahms-l] Shift report 20050425 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Mon Apr 25 2005 - 16:00:31 EDT
Shifter leader was SB
Shifters were 


1. Beam from previous night shift continued throughout the day. RHIC went to machine development around 2:20 pm
2. FS@2.3 deg @1/1A
   trig2(182k), trig3(574k), trig4(128k), trig5(225k)
3. No major detector problems. Bottom channels 13 and 14 of H2 were incorrectly cabled. Flemming put them back in their orginal configuration.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Apr 25 16:00:49 2005

This archive was generated by hypermail 2.1.8 : Mon Apr 25 2005 - 16:01:11 EDT